OpenStudy (anonymous):

Find the three arithmetic means between 12 and -16.

OpenStudy (anonymous):

5, -2, -9 5, -1, -9 6, -2, -9 6, -1, -8

mathslover (mathslover):

Oh k so do you know the formula for arithmetic means?

OpenStudy (anonymous):

no

mathslover (mathslover):

np let me tell u that one: \[\huge{\textbf{Arithmetic mean} =\frac{a+b}{2}}\] where a and b are no.

mathslover (mathslover):

so we have arithmetic mean = (12-16)/2 = -4/2 = -2 this is the first AM

OpenStudy (anonymous):

if u insert AMs the resulting five nos. will be in an Arithmetic progression now the first term is given 12 and the 5th term is -16 (as there are 3 AMs apart from 12 and -16) thus there difference is -16- 12 now the gap between the 1st term and the 5th term is of 4 times th egap between 1st and second (in AP the diffence of terms is constant) thus the constant diffeernce is -28/4=-7 thus the numbers are (12-7),(12-7-7),(12-7-7-7) i.e.5, -2, -9

mathslover (mathslover):

for finding another two we have: -16 = 12 + (28-1)d -16 = 12+ 27*d -28 = 27 d d = -28/27

OpenStudy (anonymous):

is it c?

OpenStudy (anonymous):

no its a

mathslover (mathslover):

a)

OpenStudy (anonymous):

thank you!

mathslover (mathslover):

np